OpenStudy (anonymous):

The given function f is differentiable. Verify that f has an inverse and find (f^(-1))'(c): f(x) = (x+3)/(x-1), x>1; c=3.

OpenStudy (anonymous):

Really. I'm just a little confused on how to verify that f has an inverse. Doesn't the function x have to always be increasing or decreasing in order to be one-to-one? With this function, it dips up and down (but never negative within the given parameters).

OpenStudy (anonymous):

i think this is one to one, lets check by finding the inverse

OpenStudy (anonymous):

oh to answer your question, the function does not have to be increasing or decreasing to be one to one unless it is continuous this one is not

OpenStudy (anonymous):

Here is my work. I just don't know how I go about verifying the inverse.

OpenStudy (anonymous):

I did the second part of the problem (find (f^(-1))'(c) only.

OpenStudy (anonymous):

\[x=\frac{y+3}{y-1}\] \[x(y-1)=y+3\] \[xy-x=y+3\] \[xy-y=x+3\] \[y(x-1)=x+3\] \[y=\frac{x+3}{x-1}\] so this function not only has an inverse, it is its own inverse

OpenStudy (anonymous):

What are the rules in order for a function to be one-to-one? Does it always have to be increasing or decreasing, or does it just not have to be negative?

OpenStudy (anonymous):

you can use the formula you used, but you can also take the derivative and replace \(x\) by \(3\)

OpenStudy (anonymous):

for a function \(f\) to be one to one it means if \(f(a)=f(b)\) then \(a=b\)

OpenStudy (anonymous):

you can prove the function is one to one by a) checking to see whether it passes the horizontal line test if you have the graph b) putting \(f(a)=f(b)\) and showing by algebra that \(a=b\) c) finding an inverse that is a function since you can find the inverse, it must be one to one

OpenStudy (anonymous):

Ah, ok then. And what you did was step c then, right?

OpenStudy (anonymous):

yes, since now that you have the inverse, you can clearly find its derivative

OpenStudy (anonymous):

What kind of problem would you have run into if you COULDN'T find the inverse?

OpenStudy (anonymous):

Would you just not be able to solve for y? And that's how you'd know that there wasn't an inverse?

OpenStudy (anonymous):

well you might get something like \(y=x\pm\sqrt{4-x^2}\) for example

OpenStudy (anonymous):

Ohh ok, since there would be more than one point on the graph, it wouldn't be one-to-one.

OpenStudy (anonymous):

or for a simpler example here is a simple not one to one function, \(f(x)=x^2\) if you try to solve \(y=x^2\) for \(x\) you get \(x=\pm\sqrt{y}\) so not a function

OpenStudy (anonymous):

Gotcha, thanks.